Scenario 1: Mr. Tiny was seen by Dr. Kirt, Aetna was billed out for the services; we accept assignment. Mr. Tiny has a coinsurance of 90/10; assume there is no deductible at this point or copayment.  Aetna responded as follows:

Office Visit     Charges: 190.00    Allowed: 135.00
X-ray              Charges: 35.00      Allowed: 18.75
Inj. Morphine    Charges: 50.00      Allowed: 25.00

How much would you expect Aetna to pay for his services? $_____.___ (use this format please)

Note: You will need to round down as the insurance company will not pay a penny more than they have to.

Scenario 2: Mr. Tiny was seen by Dr. Kirt, Aetna was billed out for the services; we accept assignment. Mr. Tiny has a coinsurance of 90/10; assume there is no deductible at this point or copayment. Aetna responded as follows:

Office Visit     Charges: 190.00    Allowed: 135.00
X-ray              Charges: 35.00      Allowed: 18.75
Inj. Morphine    Charges: 50.00      Allowed: 25.00

Would we be doing a write off (adjustment) for this patient? If so, how much? $_____.___ (use this format please)

Scenario 3: Mr. Tiny was seen by Dr. Kirt, Aetna was billed out for the services; we accept assignment. Mr. Tiny has a coinsurance of 90/10; assume there is no deductible at this point or copayment. Aetna responded as follows:

Office Visit     Charges: 190.00    Allowed: 135.00
X-ray              Charges: 35.00      Allowed: 18.75
Inj. Morphine    Charges: 50.00      Allowed: 25.00

Would we be billing the patient for any of his services? If so, how much? $_____.___ (use this format please)

Note: You will need to round up for this question.

Scenario 4: Mrs. Jones was seen by Dr. Kirt, Blue Cross Blue Shield was billed out for the services; we accept assignment. Mrs. Jones has a coinsurance of 70/30; assume there is no deductible at this point or copayment. Blue Cross responded as follows:

Office Visit     Charges: 145.00    Allowed: 105.00
Wart Removal Charges: 75.00      Allowed: 59.15


How much would you expect Blue Cross to pay for her services? $_____.___ (use this format please)

Note: You will need to round down as the insurance company will not pay a penny more than they have to.

Scenario 5: Mrs. Jones was seen by Dr. Kirt, Blue Cross Blue Shield was billed out for the services; we accept assignment. Mrs. Jones has a coinsurance of 70/30; assume there is no deductible at this point or copayment. Blue Cross responded as follows:

Office Visit     Charges: 145.00    Allowed: 105.00
Wart Removal Charges: 75.00      Allowed: 59.15


Would you be doing a write off (adjustment) for this patient? $_____.___ (use this format please)

Scenario 6: Mrs. Jones was seen by Dr. Kirt, Blue Cross Blue Shield was billed out for the services; we accept assignment. Mrs. Jones has a coinsurance of 70/30; assume there is no deductible at this point or copayment. Blue Cross responded as follows:

Office Visit     Charges: 145.00    Allowed: 105.00
Wart Removal Charges: 75.00      Allowed: 59.15


Would we be billing the patient for any of her charges? If so, how much? $_____.___ (use this format please)

Note: You will need to round up.
Answer #1

First of all i would like to explain about  coinsurance, deductible and copayment

Deductible

A deductible is usually a fix dollar amount that you have to pay out of your own pocket before the insurance will cover the remaining eligible expenses.Depending on the insurance plan, the deductible can range from $0 all the way up to thousands of dollars. It can also be paid per sickness/injury (per condition) or per certificate period. As a rule of thumb, the higher the deductible the lower the premium (price to buy the plan), and vice versa – but always be sure to choose the deductible that is appropriate for your circumstances when purchasing an insurance policy.

Coinsurance

Coinsurance is usually a percentage, and represents the percentage cost that you will need to pay and the insurance plan will pay towards your eligible medical expenses. Some common coinsurance examples include: 100%, 80/20, 90/10 and 50/50 – so if you have 80/20 coinsurance on your insurance plan, it means that the insurance company will cover 80% of your medical cost and you are responsible for paying the other 20% yourself. A deductible is commonly use together with coinsurance. In this case you would pay the deductible amount first and after you would have the left over coinsurance amount.

Copays

Copays are similar to deductibles, in that it is usually a fixed amount of money you have to pay each time you need to use your insurance plan. Unlike deductibles, copays tend to be smaller dollar amounts and are applied on a per visit basis so that that you would have to pay it each visit.

Scenario 1

Office visit: Charges: 190.00    Allowed: 135.00 Aetna share 90% of allowed: 121.5

X-ray: Charges: 35.00      Allowed: 18.75 Aetna share 90% of allowed: 16.875

Inj.morphine:Charges: 50.00      Allowed: 25.00   Aetna share 90% of allowed: 22.5

For example office visit

Allowed 135$ = 135*90/100 = 121.5

Aetna billed out = 121.5+16.875+ 22.5 = 160.875$

...................................................................................................................................................................

Scenario 3

Office visit:   Charges: 190.00    Allowed: 135.00 Aetna share 90% of allowed: 121.5

X-ray:   Charges: 35.00      Allowed: 18.75    Aetna share 90% of allowed: 16.875

Inj.morphine:Charges: 50.00      Allowed: 25.00   Aetna share 90% of allowed: 22.5

The patient should pay : Charges - allowed + 10% of allowed

office Visit: 190-135 + 13.5 = 55 +13.5 = 68.5

X-ray : 16.25+1.875 =18.125

inj.morphine: 25+2.5 =27.5

total 68.5+18.125+27.5 = 114.125$

...................................................................................................................................................................

Scenario-4

Office Visit     Charges: 145.00 Allowed: 105.00 Blue cross blue shield share 70% from allowed 73.5
Wart Removal Charges: 75.00 Allowed: 59.15 Blue cross blue shield share 70% from allowed 53.235

Blue cross blue shield billed out 73.5+53.235 = 53.235

.................................................................................................................................................................sScenorio-6

Office Visit     Charges: 145.00 Allowed: 105.00 Blue cross blue shield share 70% from allowed 73.5
Wart Removal Charges: 75.00 Allowed: 59.15 Blue cross blue shield share 70% from allowed 53.235

The patient should pay : Charges - allowed +30%of allowed

145- 105+31.5 = 71.5 Office Visit

75-59.15+17.745 = 33.595 Wart Removal Charges

total = 105.095 $ the patient will pay